# SIGNALS on the mixing of oxygen and nitrogen in the spiral galaxy NGC 6946

### v1

@@ Abstract
The existence of radial gradients in the present-day metallicity of spiral galaxies has been known and well-studied for decades from observations of their H ii regions (Searle 1971; Vila-Costas & Edmunds 1992; Zaritsky et al. 1994; Moustakas et al. 2010; Sánchez et al. 2014; Zurita et al. 2021) and massive stars (Urbaneja et al. 2005a, b; Kudritzki et al. 2008, 2012; Bresolin et al. 2009; Bresolin et al. 2022). On the other hand, gathering evidence for second-order metallicity variations within galaxy discs along the azimuthal coordinate has proven to be elusive using traditional long-slit spectroscopy (Kennicutt & Garnett 1996; Martin & Belley 1996; Bresolin 2011; Li et al. 2013) or other techniques (Cedrés & Cepa 2002; Rosales-Ortega et al. 2011). The order-of-magnitude increase in the number of spectra collected within individual galaxies enabled by integral field spectroscopy has led to the unequivocal detection of azimuthal oxygen abundance gradients in spiral galaxies, in the form of differences on the order of 0.05-0.1 dex between arm and interarm regions, or across the widths of spiral arms (Sánchez et al. 2015; Vogt et al. 2017; Sánchez-Menguiano et al. 2017; Ho et al. 2017, 2018). Classical chemical evolution models incorporating the effects of spiral arms on the distribution of metals, and oxygen in particular, have been introduced by Spitoni et al. (2019) and Mollá et al. (2019), who identified azimuthal fluctuations with amplitudes on the order of 0.1 dex, compatible in magnitude with those observed. Comparable results have been obtained from numerical simulations (Solar et al. 2020; Khoperskov et al. 2023).
@@
Azimuthal variations in gas-phase metallicities are, however, not universally detected (Zinchenko et al. 2016; Williams et al. 2022; Chen et al. 2024), and even when they are found, the association with the spiral arms or the galactic environment is not always manifest, with different authors drawing contrasting conclusions. For example, only half of the sample of spiral galaxies analyzed by Kreckel et al. (2019) displays azimuthal variations of the H ii region metallicities, the link with the spiral arms sometimes missing, and thus providing no definitive indication for a systematic arm vs. interarm distinction. Sánchez-Menguiano et al. (2020) instead conclude that the arms are more metal-rich than the interarm regions in about half of their sample, an effect that is reinforced in more massive and grand-design galaxies. Also Chen et al. (2024) find that metallicity variations across arms, when detected, are larger in galaxies with more prominent spiral structures, while Grasha et al. (2022) cannot establish a connection between azimuthal variations and spiral patterns. Therefore, observationally it is still unclear whether the local physical conditions of the interstellar medium (ISM) or the existence of spiral arm structures represent the main drivers for the azimuthal distribution of the gas.
@@
This paper is organized as follows: we present the observations and the catalogue of H ii regions in NGC 6946 in Section 2, and explain the methodology followed to measure the chemical abundances in Section 3. We look in detail at the azimuthal trends we derive across the NE arm for both the oxygen and nitrogen abundances in Section 4. In Section 5 we use the two-point correlation to examine the scales of homogeneity in the 2D distribution of the two elements. We summarize our results in Section 6 and present selected results from an alternative abundance diagnostic in the Appendix.
@@ 2.2 Classification and selection of integrated regions
The line fluxes were corrected for the effect of dust attenuation using the extinction curve by Cardelli et al. (1989), assuming an electron temperature = 104 K. We then used classical BPT (Baldwin et al. 1981) diagrams to remove regions that are inconsistent with photoionization by hot stars, adopting the diagnostic curves of Kauffmann et al. (2003) in the log([N ii]6583/H) vs. log([O iii]5007/H) plane and Kewley et al. (2006) in the log([S ii]6717,6731/H) vs. log([O iii]5007/H) plane.
@@ 3 Chemical abundances
001 | 308.88342 | 60.12317 | 117.7 34.0 | 133.7 6.2 | 95.5 5.5 | 477.6 9.3 | 76.1 2.0 | 52.7 6.6 | 35.8 6.6 |
002 | 308.87651 | 60.18765 | 189.5 48.6 | 94.2 9.9 | 53.2 8.1 | 355.7 10.0 | 74.8 2.2 | 43.0 7.1 | 35.9 7.2 |
003 | 308.86973 | 60.15038 | 331.4 41.5 | 234.7 10.6 | 361.3 12.3 | 924.3 13.3 | 153.9 2.9 | 89.5 9.4 | 60.7 9.4 |
004 | 308.86560 | 60.16170 | 118.0 26.3 | 143.3 9.4 | 399.6 12.7 | 498.6 6.2 | 99.1 2.0 | 77.0 6.5 | 54.8 6.5 |
005 | 308.86354 | 60.19744 | 105.3 30.1 | 65.6 6.0 | 53.5 5.5 | 216.7 7.1 | 55.1 1.6 | 40.6 5.2 | 27.3 5.2 |
006 | 308.86352 | 60.16682 | 206.4 29.0 | 134.4 5.4 | 79.6 5.5 | 632.3 12.5 | 148.2 2.8 | 103.6 9.0 | 60.3 9.0 |
007 | 308.86147 | 60.17031 | 305.5 50.3 | 157.5 8.8 | 190.3 9.3 | 683.0 14.8 | 139.3 3.3 | 120.1 10.8 | 82.5 10.7 |
008 | 308.85979 | 60.16645 | 207.6 34.6 | 129.3 9.2 | 127.3 9.0 | 696.2 13.1 | 164.8 3.0 | 129.1 9.5 | 84.6 9.5 |
009 | 308.85975 | 60.16807 | 353.7 35.1 | 228.8 9.5 | 226.2 9.3 | 1030.4 19.7 | 238.5 4.5 | 194.6 14.4 | 126.0 14.3 |
010 | 308.85973 | 60.16879 | 236.8 29.5 | 149.8 6.4 | 147.2 6.3 | 643.4 12.7 | 141.5 2.9 | 112.3 9.2 | 72.3 9.2 |
Garner et al. (2025) demonstrated that the approach we follow yields oxygen and nitrogen abundances that are compatible with those from electron temperature measurements obtained as part of the CHAOS project (Berg et al. 2015, 2020; Croxall et al. 2015, 2016). The comparison still suffers from the well-known issues that appear when comparing photoionization models to electron temperature methods (e.g., Kewley & Ellison 2008; Moustakas et al. 2010), but there is broad agreement between the two techniques. See Fig. 6 in Garner et al. (2025) and the accompanying discussion for more details.
@@ 3.1 Radial gradients
The exact functional form of these gradients is of secondary importance for the remainder of this work, which only takes abundance residuals relative to the radial trends into consideration. A potential flattening at large galactocentric distances cannot be ruled out (however, the number of data points is small), but since our analysis will be focused on the region we do not attempt to account for it. We note that the rms scatter measured about the linear regressions is 0.10 dex (for O/H) and 0.14 dex (for N/H), considerably larger than the values one measures adopting strong-line abundance diagnostics (see discussion in Kreckel et al. 2020), but comparable to those obtained through the direct method (e.g. Croxall et al. 2015; Berg et al. 2015). Lara-López et al. (2023) derived the oxygen abundance gradient in NGC 6946 from a large number of fiber spectra, adopting the S calibration by Pilyugin & Grebel (2016), and measured a scatter of 0.04 dex. Using the same technique we obtain an even smaller value, 0.03 dex. Vale Asari et al. (2016) illustrated how strong-line diagnostics can yield artificially small abundance dispersions. It is therefore important to be critical of the results delivered by such methods. Still, the fact that we obtain a larger scatter, more in line with the direct abundance method, does not imply that our Bayesian approach is necessarily producing more realistic results.
@@ 5 Homogeneity scales of oxygen and nitrogen
In this section we compare for the first time the two-point correlation of the gas-phase abundances of both oxygen and nitrogen in the same galaxy, and follow, for simplicity, the approach by Kreckel et al. (2020) to quantify the mixing scale. As in Kreckel et al. (2020) our results refer to the distribution of metals in discrete H ii regions, rather than single spaxels as in Li et al. (2021, 2023). Williams et al. (2022) followed both approaches, and found moderately consistent results. We share with the investigations based on PHANGS-MUSE data also the excellent physical spatial resolution: 30 pc in our case (40-70 pc in Kreckel et al. 2020), compared to the hundreds of parsecs average resolution of the data analysed by Li et al. (2021, 2023).
@@
where and are the vector positions of the H ii regions where the metal field is measured, with , the angle brackets indicate the average over all possible choices of , and the horizontal bars the average over all H ii regions. In our case refers to the abundance fluctuations of either O/H or N/H. According to KT18, Eq. 3 can be simplified to
where
is the variance of . For the evaluation of the two-point correlation from Eq. 4 we calculate the product for each pair of H ii regions () separated by the distance , and divide the full sample into 0.1 kpc-wide bins in within the range 0-3 kpc.222We adapted the code by Williams et al. (2022), available at https://github.com/thomaswilliamsastro/metallicity_gpr/tree/master. By construction, the correlation decreases from a maximum value (each H ii region correlating perfectly with itself) to zero at very large separations. We execute 100 realizations by Gaussian sampling the O/H and N/H residuals uncertainties in order to estimate the 1 widths of the correlation curves. In addition, we run 100 realizations where we randomize the abundance residuals distribution in order to derive the curve corresponding to a lack of correlation between neighbouring H ii regions.
The results are shown in Fig. 9, where it is evident that the curves (actually bands) for both oxygen and nitrogen deviate significantly from the curve obtained by randomizing the abundances, and that the N/H correlation curve lies consistently below the O/H curve. To estimate the significance of the offset, we randomly generated curve pairs (one curve for O/H and one for N/H) lying within the respective lower and upper bounds, each time calculating their offset along the -axis, performing the KS test and deriving a -value from the test. Since the curves touch each other at , we avoided separations kpc. We find that the mean -value is 0.026 with a standard deviation of 0.019. This implies that the offset seen in Fig. 9 can be considered as statistically significant, i.e. the nitrogen correlation is slightly smaller than that of oxygen.
@@
The correlation scales that we measure fall within the range derived (for oxygen) for other nearby galaxies (see comparison in Li et al. 2021). The values for O/H and N/H differ only at the 1.5 level, which is less significant than that of the overall offset between the curves, because these scales are measured on the steep, descending portions of the curves themselves. The difference is only appreciated at large separations. Although the trend goes in the direction expected by KT18, we do not find evidence for N/H to be characterized by a much smaller correlation ( pc at the 50 per cent level) compared to O/H, as indicated by the KT18 model with the adopted parameters (see their Fig. 4). Our result suggests that adjustments to the KT18 model parameters are called for, in particular with regard to the small injection width of AGB stars. We argue that this is justified by the significantly longer timescale over which AGB stars release their nucleosynthetic products into the ISM compared to massive stars. In a single stellar population, such a timescale is on the order of 250 Myr, over which migration of the stars is likely to take place (Johnson et al. 2023), thus leading to a large spatial extent over which nitrogen can be injected into the ISM.

### v2

@@ Abstract
The existence of radial gradients in the present-day metallicity of spiral galaxies has been known and well-studied for decades from observations of their H ii regions (e.g. Searle 1971; Vila-Costas & Edmunds 1992; Zaritsky et al. 1994; Moustakas et al. 2010; Sánchez et al. 2014; Zurita et al. 2021) and massive stars (Urbaneja et al. 2005a, b; Kudritzki et al. 2008, 2012; Bresolin et al. 2009; Bresolin et al. 2022). On the other hand, gathering evidence for second-order metallicity variations within galaxy discs along the azimuthal coordinate has proven to be elusive using traditional long-slit spectroscopy (Kennicutt & Garnett 1996; Martin & Belley 1996; Bresolin 2011; Li et al. 2013) or other techniques (Cedrés & Cepa 2002; Rosales-Ortega et al. 2011). The order-of-magnitude increase in the number of spectra collected within individual galaxies enabled by integral field spectroscopy has led to the unequivocal detection of azimuthal oxygen abundance gradients in spiral galaxies, in the form of differences on the order of 0.05-0.1 dex between arm and interarm regions, or across the widths of spiral arms (Sánchez et al. 2015; Vogt et al. 2017; Sánchez-Menguiano et al. 2017; Ho et al. 2017, 2018). Classical chemical evolution models incorporating the effects of spiral arms on the distribution of metals, and oxygen in particular, have been introduced by Spitoni et al. (2019, see also ) and Mollá et al. (2019), who identified azimuthal fluctuations with amplitudes on the order of 0.1 dex, compatible in magnitude with those observed. Comparable results have been obtained from numerical simulations (Solar et al. 2020; Khoperskov et al. 2023).
@@
Azimuthal variations in gas-phase metallicities are, however, not universally detected (Zinchenko et al. 2016; Williams et al. 2022; Chen et al. 2024), and even when they are found, the association with the spiral arms or the galactic environment is not always manifest, with different authors drawing contrasting conclusions. For example, only half of the sample of spiral galaxies analyzed by Kreckel et al. (2019) displays azimuthal variations of the H ii region metallicities, the link with the spiral arms sometimes missing, and thus providing no definitive indication for a systematic arm vs. interarm distinction. Sánchez-Menguiano et al. (2020) instead conclude that the arms are more metal-rich than the interarm regions in about half of their sample, an effect that is reinforced in more massive and grand-design galaxies. Also Chen et al. (2024) find that metallicity variations across arms, when detected, are larger in galaxies with more prominent spiral structures, while Grasha et al. (2022) cannot establish a connection between azimuthal variations and spiral patterns. A correlation between spiral arms and metal enrichment has also been detected in the Milky Way from samples of giant stars observed by Gaia (Poggio et al. 2022; Hawkins 2023; Barbillon et al. 2025). We can draw the conclusion that observationally it is still unclear whether the local physical conditions of the interstellar medium (ISM) or the existence of spiral arm structures represent the main drivers for the azimuthal distribution of the gas.
@@
This paper is organized as follows: we present the observations and the catalogue of H ii regions in NGC 6946 in Section 2, and explain the methodology followed to measure the chemical abundances in Section 3. We look in detail at the azimuthal trends we derive across the NE arm for both the oxygen and nitrogen abundances in Section 4. In Section 5 we use the two-point correlation function and an additional statistical tool, the semivariogram, to examine the scales of homogeneity in the 2D distribution of the two elements. We summarize our results in Section 6 and present selected results from an alternative abundance diagnostic in the Appendix.
@@ 2.2 Classification and selection of integrated regions
The line fluxes were corrected for the effect of dust extinction using the curve by Cardelli et al. (1989), assuming an electron temperature = 104 K. We then used classical BPT (Baldwin et al. 1981) diagrams to remove regions that are inconsistent with photoionization by hot stars, adopting the diagnostic curves of Kauffmann et al. (2003) in the log([N ii]6583/H) vs. log([O iii]5007/H) plane and Kewley et al. (2006) in the log([S ii]6717,6731/H) vs. log([O iii]5007/H) plane.
@@ 3 Chemical abundances
| 001 | 308.88342 | 60.12317 | 117.7 34.0 | 133.7 6.2 | 95.5 5.5 | 477.6 9.3 | 76.1 2.0 | 52.7 6.6 | 35.8 6.6 |
| 002 | 308.87651 | 60.18765 | 189.5 48.6 | 94.2 9.9 | 53.2 8.1 | 355.7 10.0 | 74.8 2.2 | 43.0 7.1 | 35.9 7.2 |
| 003 | 308.86973 | 60.15038 | 331.4 41.5 | 234.7 10.6 | 361.3 12.3 | 924.3 13.3 | 153.9 2.9 | 89.5 9.4 | 60.7 9.4 |
| 004 | 308.86560 | 60.16170 | 118.0 26.3 | 143.3 9.4 | 399.6 12.7 | 498.6 6.2 | 99.1 2.0 | 77.0 6.5 | 54.8 6.5 |
| 005 | 308.86354 | 60.19744 | 105.3 30.1 | 65.6 6.0 | 53.5 5.5 | 216.7 7.1 | 55.1 1.6 | 40.6 5.2 | 27.3 5.2 |
| 006 | 308.86352 | 60.16682 | 206.4 29.0 | 134.4 5.4 | 79.6 5.5 | 632.3 12.5 | 148.2 2.8 | 103.6 9.0 | 60.3 9.0 |
| 007 | 308.86147 | 60.17031 | 305.5 50.3 | 157.5 8.8 | 190.3 9.3 | 683.0 14.8 | 139.3 3.3 | 120.1 10.8 | 82.5 10.7 |
| 008 | 308.85979 | 60.16645 | 207.6 34.6 | 129.3 9.2 | 127.3 9.0 | 696.2 13.1 | 164.8 3.0 | 129.1 9.5 | 84.6 9.5 |
| 009 | 308.85975 | 60.16807 | 353.7 35.1 | 228.8 9.5 | 226.2 9.3 | 1030.4 19.7 | 238.5 4.5 | 194.6 14.4 | 126.0 14.3 |
| 010 | 308.85973 | 60.16879 | 236.8 29.5 | 149.8 6.4 | 147.2 6.3 | 643.4 12.7 | 141.5 2.9 | 112.3 9.2 | 72.3 9.2 |
Moustakas et al. (2010) found that empirical calibrations can underestimate the “true" -based gas-phase metallicity by 0.2-0.3 dex, while theoretical calibrations, based on photoionization models, yield abundances that are too high by the same amount. Consequently, the absolute uncertainty in the nebular abundance scale is about 0.7 dex (Kewley & Ellison 2008). Garner et al. (2025) demonstrated that the approach we follow yields oxygen and nitrogen abundances that are compatible with these systematic uncertainties when comparing to those made with electron temperature measurements obtained as part of the CHAOS project (Berg et al. 2015, 2020; Croxall et al. 2015, 2016).
@@ 3.1 Radial gradients
The exact functional form of these gradients is of secondary importance for the remainder of this work, which only takes abundance residuals relative to the radial trends into consideration. Because of the systematic differences affecting the nebular abundance diagnostics, mentioned above, the linear regression parameters and rms scatter about the fit can depend on the chosen abundance determination method, but our main conclusions are robust. A potential flattening at large galactocentric distances cannot be ruled out (however, the number of data points is small), but since our analysis will be focused on the region we do not attempt to account for it. We note that the rms scatter measured about the linear regressions is 0.10 dex (for O/H) and 0.14 dex (for N/H), considerably larger than the values one measures adopting strong-line abundance diagnostics (see discussion in Kreckel et al. 2020), but comparable to those obtained through the direct method (e.g. Croxall et al. 2015; Berg et al. 2015). Lara-López et al. (2023) derived the oxygen abundance gradient in NGC 6946 from a large number of fiber spectra, adopting the S calibration by Pilyugin & Grebel (2016), and measured a scatter of 0.04 dex. Using the same technique we obtain an even smaller value, 0.03 dex.
@@ 5 Homogeneity scales of oxygen and nitrogen
5.1 Two-point correlation function
In this section we compare for the first time the two-point correlation of the gas-phase abundances of both oxygen and nitrogen in the same galaxy, and follow, for simplicity, the approach by Kreckel et al. (2020) to quantify the mixing scale. As in Kreckel et al. (2020), our results refer to the distribution of metals in discrete H ii regions, rather than single spaxels as in Li et al. (2021, 2023). Williams et al. (2022) followed both approaches, and found moderately consistent results. We share with the investigations based on PHANGS-MUSE data also the excellent physical spatial resolution: 30 pc in our case (40-70 pc in Kreckel et al. 2020), compared to the hundreds of parsecs average resolution of the data analysed by Li et al. (2021, 2023).
@@ 5.1 Two-point correlation function
where and are the vector positions of the H ii regions where the metal field is measured, the angle brackets indicate the average over all possible choices of , and the horizontal bars represent the average over all H ii regions. In our case refers to the logarithmic abundance fluctuations of either O/H or N/H, with a mean value of zero. According to KT18 (see also Li et al. 2021), Eq. 3 can then be simplified to
where is the variance of . For the evaluation of the two-point correlation from Eq. 4 we calculate the product for each pair of H ii regions () separated by the distance , and divide the full sample into 0.1 kpc-wide bins in within the range 0-4 kpc.222We adapted the code by Williams et al. (2022), available at https://github.com/thomaswilliamsastro/metallicity_gpr/tree/master. By construction, the correlation decreases from a maximum value (each H ii region correlating perfectly with itself) to zero at very large separations. We execute 100 realizations by Gaussian sampling the log(O/H) and log(N/H) residuals uncertainties in order to estimate the 1 widths of the correlation curves. In addition, we run 100 realizations where we randomize the abundance residuals distribution in order to derive the curve corresponding to a lack of correlation between neighbouring H ii regions.
Two distinct approaches have been followed in the recent literature concerning the separation between pairs in Eq. 3 and 4: Kreckel et al. (2020) and Williams et al. (2022) adopt , while KT18 and Li et al. (2021, 2023) utilize . The adoption of the equality sign in the latter case is in line with other uses of the two-point correlation function in astronomy (e.g. Peebles 1993). We utilize both definitions below, and show that they provide compatible outcomes.
(a) Our results adopting are displayed in Fig. 9a, where it is evident that the curves (actually bands) for both oxygen and nitrogen deviate significantly from the curve obtained by randomizing the abundances, and that the N/H correlation curve lies consistently below the O/H curve. To estimate the significance of the offset, we randomly generated curve pairs (one curve for O/H and one for N/H) lying within the respective lower and upper bounds, each time calculating their offset along the -axis, performing the KS test and deriving a -value from the test. Since the curves touch each other at , we avoided separations kpc. We find that the mean -value is 0.026 with a standard deviation of 0.019. This implies that the offset seen in Fig. 9 can be considered as statistically significant, i.e. the nitrogen correlation is slightly smaller than that of oxygen.
@@
The correlation scales that we measure fall within the range derived (in the case of oxygen) for other nearby galaxies (see comparison in Li et al. 2021). The values for O/H and N/H differ only at the 1.5 level, which is less significant than that of the overall offset between the curves, because these scales are measured on the steep, descending portions of the curves themselves. The difference is only appreciated at large separations.
(b) If we use the standard definition of the two-point correlation, where , we obtain the curves shown in Fig. 9b, which look different from those illustrated above. In particular, they drop sharply at small separations and display significant fluctuations in amplitude, due to the stricter criterion that isolates a much smaller number of H II regions per bin. The correlation function values are also considerably lower than in the previous case – such low values are compatible with those calculated by Li et al. (2021, see their Fig. 3). Nevertheless, the location of the N/H curve below the O/H curve is retained.
@@
Although the trend we see in Fig. 9 goes in the direction expected by KT18, we do not find evidence for N/H to be characterized by a much smaller correlation scale ( 100 pc at the 50 per cent level) compared to O/H, as indicated by the KT18 model with the adopted parameters (see their Fig. 4). Our result suggests that adjustments to the KT18 model parameters are called for, in particular with regard to the small injection width of AGB stars. We argue that this is justified by the significantly longer timescale over which AGB stars release their nucleosynthetic products into the ISM, in comparison to massive stars. In a single stellar population, such a timescale is on the order of 250 Myr, over which migration of the stars is likely to take place (Johnson et al. 2023), thus leading to a large spatial extent over which nitrogen can be injected into the ISM.
5.2 Semivariogram
In order to consolidate our conclusions, we also explored the use of the semivariogram, a geostatistical tool that has been recently applied to the analysis of the 2D chemical abundance distributions in galaxies by Metha et al. (2021); Metha et al. (2022). In our case the semivariogram measures how the variance of the chemical abundances changes as a function of separation between pairs of H II regions. Like the two-point correlation function, which is a strictly associated concept, it can inform us on the spatial scales of homeogeneity of the chemical abundance fields. Using the quantities defined in Eq. 3:
where the variance is calculated over all pairs of H II regions () separated by a distance such that , with the chosen bin size (Metha et al. 2021).
The semivariograms for both log(O/H) and log(N/H), adopting kpc, are shown in Fig. 10. The vertical offset between the two is due to the larger (uncorrelated) measurement uncertainties of the N/H abundances, since the value increases with the measurement errors (Metha et al. 2021). In our sample, the mean errors are 0.05 dex for log(O/H) and 0.07 dex for log(N/H). The observed semivariances at are considerably larger (by a factor of 3.5-4) than what is derived from the mean errors, suggesting the presence of correlated abundance fluctuations on smaller spatial scales.
The amplitude of the chemical abundance fluctuations can be estimated by the difference in measured between large separations, where the abundances are uncorrelated and the curve flattens out, and small separations. Such a difference is 0.006-0.008 dex, implying fluctuations on the order of 0.11-0.12 dex.
In order to estimate the spatial scales over which the chemical abundances can be considered to be correlated, we followed Metha et al. (2022) and fitted the semivariograms with a simple model:
where is the offset measured at , corresponds to the value described above, and is a measure of the homogeneity scale of the chemical abundances. Our fits are shown as dashed lines in Fig. 10. We estimated the errors in the fit parameters by varying the range in over which the fit is performed (5-20 kpc in 0.1 kpc steps). We obtain kpc for log(O/H) and kpc for log(N/H). The marginal difference is consistent with our previous finding (Sect. 5.1) that the correlation scale of N/H is slightly smaller than that of O/H. We calculate that the correlation functions for O/H and N/H shown in Fig. 9a fall to 33 per cent at a separation .
In summary, using different statistical tools, we find evidence that chemical abundance fluctuations on the order of 0.1 dex exist in NGC 6946 for both oxygen and nitrogen. Although nitrogen is slightly less correlated than oxygen, the spatial correlation is comparable between the two elements, and is on the order of half a kpc.